Find the cube root of each of the following numbers by prime factorisation method.

 

(vi) 13824

 

 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

(vi) 13824

Solution:

13824 = 2×2×2×2×2×2×2×2×2×3×3×3

By grouping the factors in triplets of equal factors,

13824 = (2×2×2)×(2×2×2)×(2×2×2)×(3×3×3)

Here, 13824 can be grouped into triplets of equal factors,

∴ 13824 = (2×2× 2×3) = 24

Hence, 24 is cube root of 13824.




Thanks